# Stability and Generalization for Bellman Residuals

### v1

@@ 1 Introduction
Modern decision-making systems—from sepsis treatment strategies in intensive-care units to route planning for autonomous vehicles—must reason about sequences of actions whose consequences only unfold over time. Reinforcement learning (RL) provides a principled framework for such dynamic problems, formalizing them as Markov decision processes (MDPs) and prescribing policies that optimize long-horizon rewards. Yet in many high-stakes domains on-line interaction is either unethical, dangerous, or simply too expensive Jiang and Xie (2024). Practitioners therefore turn to offline RL, where the learner is handed a fixed batch of interaction data collected by some behaviour policy, and to its sister field inverse RL (IRL), which infers underlying preferences (i.e., reward functions) from logged expert trajectories. Across both settings, the crux of learning remains the same: estimate a value function that satisfies the Bellman optimality equations even though no new state–action pairs can be queried.
Unfortunately, enforcing Bellman consistency offline is notoriously difficult. Policy-gradient objectives (Mei et al., 2020; Cen et al., 2022) assume online sampling and misalign with a static dataset. Fitted fixed-point methods, exemplified by Fitted Q-Iteration (Ernst et al., 2005) or Fitted Value Iteration (Munos and Szepesvári, 2008), can diverge catastrophically once the “deadly triad” of function approximation, bootstrapping, and off-policy data is present (Tsitsiklis and Van Roy, 1996; Van Hasselt et al., 2018). Importance-weighting approaches mitigate distribution shift but leave no guarantee that the learned $Q$ satisfies the Bellman equations uniformly over all $(s,a)$ pairs (Jiang and Xie, 2024). Bellman residual minimisation (BRM), directly fitting $Q$ by driving the squared Bellman error to zero, has long been viewed as conceptually appealing and practically effective, yet with limited theory to back global optimality convergence (Jiang and Xie, 2024).
@@
Section 2 formalizes the discounted MDP setting and shows how Bellman-residual minimization can be cast as a minimax problem. It also details the Stochastic Gradient Descent–Ascent (SGDA) algorithm and establishes its global convergence under Polyak–Łojasiewicz geometry. Section 3 develops our main statistical results: an $\mathcal{O}(1/n)$ on-average argument-stability bound for SGDA and the resulting $\mathcal{O}(1/n)$ generalization and excess-risk guarantees for Bellman residual minimization. All technical proofs and auxiliary lemmas are collected in the Appendix.
@@ 2.2 Bellman Residual Minimization
A canonical way of solving the mini-max problem is to apply the Stochastic Gradient Ascent Descent (SGDA) algorithm (Yang et al., 2020). Kang et al. (2025) proved that both Equation (3) and (4) satisfy the Polyak-Łojasiewicz condition for the parametrization with popular function classes such as Neural Network, and therefore the SGDA algorithm finds the $(\boldsymbol{\theta}_{1},\boldsymbol{\theta}_{2})$ that are global minima of Equation (4). In the following Section 2.3, we elaborate on the SGDA algorithm.
@@ 4 $g_{t}^{v}\leftarrow\frac{1}{B}\sum_{i\in I_{t}}\nabla_{v}f(w_{t},v_{t};z_{i})$;
Let the iterates in Algorithm 1 be written as $\{(\boldsymbol{\theta}_{2,t},\boldsymbol{\theta}_{1,t})\}_{t\geq 0}$, where $\boldsymbol{\theta}_{2,t}$ parametrises the action–value function $Q_{\boldsymbol{\theta}_{2,t}}$ (primal variable) and $\boldsymbol{\theta}_{1,t}$ parametrises $\zeta_{\boldsymbol{\theta}_{1,t}}$ (dual variable). Choose the harmonic step-sizes $\eta_{t}=\frac{c_{1}}{\,c_{2}+t\,},t\geq 0$ with some constants $c_{1}>0$ and $c_{2}\geq 1$ such that $\eta_{t}\leq\min\!\{1/(4L),\,1/\rho\}$ for every $t$. Then the Algorithm 1’s output sequence $\{(\boldsymbol{\theta}_{2,t},\boldsymbol{\theta}_{1,t})\}$ converges almost surely to the unique saddle point of the empirical objective (4), where the suboptimality of empirical objective (4) is bounded by $\frac{d_{1}}{d_{2}+t}$ for some constants $d_{1}$ and $d_{2}$.
@@ 3.1 Stability
    QG for $\Phi_{D}$. $\Phi_{D}$ satisfies Quadratic Growth (QG) with parameter $\mu_{\mathrm{QG}}>0$: $\Phi_{D}(w)-\Phi_{D}^{\star}\geq\tfrac{\mu_{\mathrm{QG}}}{2}\|w-x_{D}^{\star}\|^{2}$.
@@
  8. (A8)
  9. (A9)
    Unique saddle point. Each dataset $D$ admits a unique saddle point $(x_{D}^{\star},v_{D}^{\star})$.
Kang et al. (2025) proved that Assumptions (A1), (A2), and (A5) hold for the problem of minimizing Equation (4). In addition, they showed that the Equation (4) is of $\mathcal{C}^{2}$ and therefore the equivalence of PL and QG holds by Liao et al. (2024), satisfying Assumption (A3).
@@ Appendix A Notations
$L$ Smoothness (Assumption A1)
$\mu_{PL}$ Polyak–Łojasiewicz condition constant (Assumption A2)
$\mu_{QG}$ Quadratic Growth constant (Assumption A3)
$\rho$ Strong concavity constant (Assumption A4)
$G$ Per-sample gradient bound (Assumption A5)
@@ B.1 Supporting Lemmas for proving Theorem 3
Let $D^{(i)}$ be obtained from $D$ by replacing one sample. Assume (A1) (joint $L$-smoothness), (A4) ($\rho$-strong concavity in $v$), (A5) (per-sample gradients bounded by $G$), (A2)(A3) (PL and QG for $\Phi_{D}$), and (A8) (the same constants hold for $D$ and $D^{(i)}$). Let $x^{\star}_{D}\in\arg\min_{w}\Phi_{D}(w)$ and $x^{\star}_{D^{(i)}}\in\arg\min_{w}\Phi_{D^{(i)}}(w)$, and define $v^{\star}_{D}:=v_{D}^{\star}(x^{\star}_{D})$ and $v^{\star}_{D^{(i)}}:=v_{D^{(i)}}^{\star}(x^{\star}_{D^{(i)}})$. Then
@@
Recall $w^{+}=w_{t}-\eta\,g^{w}$ with $g^{w}:=\nabla_{w}f(w_{t},v_{t};z_{p}hi-smooth-step{i_{t}})$ and let $\eta:=\eta_{t}$. By $L_{\Phi}$-smoothness of $\Phi$,
@@
$\leq\|w_{t}-x_{D}^{*}\|+\|x_{D}^{*}-x_{D^{(i)}}^{*}\|+\|x_{D^{(i)}}^{*}-w_{t}^{\prime}\|+\|v_{t}-v_{D}^{*}\|+\|v_{D}^{*}-v_{D^{(i)}}^{*}\|-\|v_{D^{(i)}}^{*}-v_{t}^{\prime}\|$ (Triangle inequality)
$\leq C_{\mathrm{dist}}\left(\sqrt{\Psi_{\alpha,\mathcal{D}}(w_{t},v_{t})}+\sqrt{\Psi_{\alpha,\mathcal{D}^{(i)}}(w_{t}^{\prime},v_{t}^{\prime})}\right)+\|x_{D}^{*}-x_{D^{(i)}}^{*}\|-\|v_{D^{(i)}}^{*}-v_{t}^{\prime}\|$ (Lemma 12)

### v2

@@ 1 Introduction
Modern decision‐making systems—from sepsis treatment strategies in intensive-care units to route planning for autonomous vehicles—must reason about sequences of actions whose consequences only unfold over time. Reinforcement learning (RL) provides a principled framework for such dynamic problems, formalizing them as Markov decision processes (MDPs) and prescribing policies that optimize long-horizon rewards. Yet in many high-stakes domains on-line interaction is either unethical, dangerous, or simply too expensive Jiang and Xie (2024). Practitioners therefore turn to offline RL, where the learner is handed a fixed batch of interaction data collected by some behaviour policy, and to its sister field inverse RL (IRL), which infers underlying preferences (i.e., reward functions) from logged expert trajectories. Across both settings, the crux of learning remains the same: estimate a value function that satisfies the Bellman optimality equations even though no new state–action pairs can be queried.
Unfortunately, enforcing Bellman consistency offline is notoriously difficult. Policy-gradient objectives (Mei et al., 2020; Cen et al., 2022) assume online sampling and misalign with a static dataset. Fitted fixed-point methods, exemplified by Fitted Q-Iteration (Ernst et al., 2005) or Fitted Value Iteration (Munos and Szepesvári, 2008), can diverge catastrophically once the “deadly triad” of function approximation, bootstrapping, and off-policy data is present (Tsitsiklis and Van Roy, 1996; Van Hasselt et al., 2018; Chen et al., 2023). Importance-weighting approaches mitigate distribution shift but leave no guarantee that the learned $Q$ satisfies the Bellman equations uniformly over all $(s,a)$ pairs (Jiang and Xie, 2024). Bellman residual minimisation (BRM), directly fitting $Q$ by driving the squared Bellman error to zero, has long been viewed as conceptually appealing and practically effective, yet with limited theory to back global optimality convergence (Jiang and Xie, 2024).
@@
Section 2 formalizes the discounted MDP setting and shows how Bellman‐residual minimization can be cast as a minimax problem. It also details the Stochastic Gradient Descent–Ascent (SGDA) algorithm and establishes its global convergence under Polyak–Łojasiewicz geometry. Section 3 develops our main statistical results: an $\mathcal{O}(1/n)$ on-average argument-stability bound for SGDA and the resulting $\mathcal{O}(1/n)$ generalization and excess-risk guarantees for Bellman residual minimization. All technical proofs and auxiliary lemmas are collected in the Appendix.
@@ 2.2 Bellman Residual Minimization
A canonical way of solving the mini-max problem is to apply the Stochastic Gradient Ascent Descent (SGDA) algorithm (Yang et al., 2020). Kang et al. (2025) proved that both Equation (3) and (4) satisfy the Polyak-Łojasiewicz condition for the parametrization with popular function classes such as Neural Network, and therefore SGDA converges to a global saddle point of (4). When the saddle set is not a singleton, the algorithm (together with the fixed initialization) selects a particular limit point, which we treat as the selected saddle. In the following Section 2.3, we elaborate on the SGDA algorithm.
@@ 4 $g_{t}^{v}\leftarrow\frac{1}{B}\sum_{i\in I_{t}}\nabla_{v}f(w_{t},v_{t};z_{i})$;
Let the iterates in Algorithm 1 be written as $\{(\boldsymbol{\theta}_{2,t},\boldsymbol{\theta}_{1,t})\}_{t\geq 0}$, where $\boldsymbol{\theta}_{2,t}$ parametrises the action–value function $Q_{\boldsymbol{\theta}_{2,t}}$ (primal variable) and $\boldsymbol{\theta}_{1,t}$ parametrises $\zeta_{\boldsymbol{\theta}_{1,t}}$ (dual variable). Choose the harmonic step-sizes $\eta_{t}=\frac{c_{1}}{\,c_{2}+t\,},t\geq 0$ with some constants $c_{1}>0$ and $c_{2}\geq 1$ such that $\eta_{t}\leq\min\!\{1/(4L),\,1/\rho\}$ for every $t$. Then the Algorithm 1’s output sequence $\{(\boldsymbol{\theta}_{2,t},\boldsymbol{\theta}_{1,t})\}$ converges almost surely to a saddle point of the empirical objective (4), where the suboptimality of empirical objective (4) is bounded by $\frac{d_{1}}{d_{2}+t}$ for some constants $d_{1}$ and $d_{2}$.
@@ 3.1 Stability
Let $X_{D}^{\star}:=\arg\min_{w}\Phi_{D}(w)$ denote the (possibly set-valued) minimizer set. When $X_{D}^{\star}$ is not a singleton, we fix a deterministic selection $x_{D}^{\star}\in X_{D}^{\star}$ (e.g. the almost-sure limit point of SGDA started from the common initialization $(w_{0},v_{0})$; see Assumption (A4) below), and define $v_{D}^{\star}:=\arg\max_{v}F_{D}(x_{D}^{\star},v)$, which is unique by (A7).
@@
    Bounded gradients on the effective domain. There exists a compact convex set $\Omega\subset\mathcal{W}\times\mathcal{V}$ such that the two coupled sequences of iterates $\{(w_{t},v_{t})\}_{t=0}^{T}$ and $\{(w^{\prime}_{t},v^{\prime}_{t})\}_{t=0}^{T}$ generated on $\mathcal{D}$ and $\mathcal{D}^{(i)}$ remain within $\Omega$ almost surely. We define $G$ as the uniform gradient bound on this set:
    $G:=\sup_{(w,v)\in\Omega,z\in\mathcal{Z}}\max\left\{\|\nabla_{w}f(w,v;z)\|,\|\nabla_{v}f(w,v;z)\|\right\}<\infty.$
  3. (A3)
  4. (A4)
    Selected saddle point. Each dataset $D$ admits at least one saddle point, and the SGDA iterates initialized at $(w_{0},v_{0})$ converge to an initialization-dependent saddle point $(x_{D}^{\star},v_{D}^{\star})$, which we call the selected saddle.
  5. (A5)
@@
    QG for $\Phi_{D}$ on the effective domain. Let $\Omega$ be the compact set from Assumption (A2). Then $\Phi_{D}$ satisfies Quadratic Growth (QG) around the selected minimizer $x_{D}^{\star}$ on $\Omega_{w}:=\{\,w:\exists v\text{ s.t. }(w,v)\in\Omega\,\}$:
    $\Phi_{D}(w)-\Phi_{D}^{\star}\geq\tfrac{\mu_{\mathrm{QG}}}{2}\|w-x_{D}^{\star}\|^{2}\qquad\forall\,w\in\Omega_{w}.$
@@
Kang et al. (2025) proved that Assumptions (A1), (A5), and (A7) hold for the problem of minimizing Equation (4). In addition, they showed that the Equation (4) is of $\mathcal{C}^{2}$ and therefore the equivalence of PL and QG holds by Liao et al. (2024), satisfying Assumption (A6). Assumption (A2) is justified by the coercivity of the PL-Strongly Concave landscape, which ensures iterates remain in a bounded sub-level set (Yang et al., 2020). While strong concavity implies unbounded gradients on $\mathbb{R}^{d}$, the geometry of problem (4) induces a drift that keeps iterates bounded (coercivity). Thus, we only require the gradient to be bounded within the effective domain $\Omega$ visited by the algorithm, consistent with the global convergence guarantees for unprojected SGDA in this setting (Yang et al., 2020). Yang et al. (2020) also proved that (A4) holds. The Assumption (A3) is standard in the stability literature: for example, Hardt et al. (2016) assume each per-example loss $f(\cdot;z)$ is $L$-Lipschitz and $\beta$-smooth uniformly in $z$, and Wang et al. (2022) assume the gradients and smoothness of $f(w,v;z)$ are bounded by global constants $G$ and $L$ for all $z$. These conditions immediately imply that the corresponding constants are identical for any dataset and its replace-one neighbor. Assumption (A3) is the PL/QG analogue of these standard uniform assumptions.
@@ Appendix A Notations
$L$ Smoothness (Assumption (A1))
$\mu_{PL}$ Polyak–Łojasiewicz condition constant (Assumption (A5))
$\mu_{QG}$ Quadratic Growth constant (Assumption (A6))
$\rho$ Strong concavity constant (Assumption (A7))
$G$ Per-sample gradient bound (Assumption (A2))
@@ B.1 Supporting Lemmas for proving Theorem 3
Let $D^{(i)}$ be obtained from $D$ by replacing one sample. Assume (A1) (joint $L$-smoothness), (A7) ($\rho$-strong concavity in $v$), (A2) (per-sample gradients bounded by $G$), (A5)(A6) (PL and QG for $\Phi_{D}$), and (A3) (the same constants hold for $D$ and $D^{(i)}$). Let $\left(x_{D}^{\star},v_{D}^{\star}\right)$ and $\left(x_{D^{(i)}}^{\star},v_{D^{(i)}}^{\star}\right)$ be the selected saddle points from Assumption (A4) for $D$ and $D^{(i)}$, respectively. Then
@@
Recall $w^{+}=w_{t}-\eta\,g^{w}$ with $g^{w}:=\nabla_{w}f(w_{t},v_{t};z_{i_{t}})$ and let $\eta:=\eta_{t}$. By $L_{\Phi}$-smoothness of $\Phi$,
@@
$\leq\|w_{t}-x_{D}^{*}\|+\|x_{D}^{*}-x_{D^{(i)}}^{*}\|+\|x_{D^{(i)}}^{*}-w_{t}^{\prime}\|+\|v_{t}-v_{D}^{*}\|+\|v_{D}^{*}-v_{D^{(i)}}^{*}\|+\|v_{D^{(i)}}^{*}-v_{t}^{\prime}\|$ (Triangle inequality)
$\leq C_{\mathrm{dist}}\left(\sqrt{\Psi_{\alpha,\mathcal{D}}(w_{t},v_{t})}+\sqrt{\Psi_{\alpha,\mathcal{D}^{(i)}}(w_{t}^{\prime},v_{t}^{\prime})}\right)+\|x_{D}^{*}-x_{D^{(i)}}^{*}\|+\|v_{D^{(i)}}^{*}-v_{t}^{\prime}\|$ (Lemma 12)